OpenStudy (anonymous):

What is the maximum product of two positive numbers whose sum is 78?

OpenStudy (amistre64):

sounds like a consecutive number thing

OpenStudy (anonymous):

1520?

OpenStudy (amistre64):

10+1 = 11; 10*1 = 10 6+5 = 11; 6*6 = 30

OpenStudy (amistre64):

78/2 = 35+4 = 39 ... hmm

OpenStudy (anonymous):

39^2 = 1521. 39-1 = 38, 78-38 = 40. 38*40 = 1520. My flawed reasoning anyway.

OpenStudy (anonymous):

Assuming he doesn't mean two of the same number?

OpenStudy (amistre64):

:) yeah, doesnt say the numbers cant be the same

OpenStudy (anonymous):

maximum is 39^2 using calculus

OpenStudy (anonymous):

1521

OpenStudy (anonymous):

Yeah. I for some reason read "positive number" as "even number."

OpenStudy (anonymous):

1521, using inequalities.